OpenStudy (anonymous):

I Can't stand Summer School Please Help evaluate 14c5 A.537,824 B.240,240 C.2,002 D70

OpenStudy (nurali):

\[nCr=\frac{ n! }{ r!(n-r)! }\] \[14C5=\frac{ 14! }{ 5!(14-5)!}\] \[\frac{ 14! }{ 5!*9! }\] \[\frac{ 14*13*12*11*10*9! }{ 5!*9! }\] \[\frac{ 240240 }{ 120 }\] \[2002\]

OpenStudy (anonymous):

Thanks, This Summer School work is very Hard

OpenStudy (anonymous):

@Nurali

OpenStudy (nurali):

Anytime.